A manager claims that the company has given an average salary increment of RM 40

to all the employees per year. A researcher wants to check if this claim is true. A

random sample of 45 employees taken by the researcher showed that these employees

got an average increment of RM 38 per year with a standard deviation of RM 8.

a) Is there enough evidence to accept the manager’s claim at 0.1 level of significance?

b) Test at 2.5% significance level whether the average salary increment of all

employees is less than RM 40 per year.


1
Expert's answer
2021-02-24T06:00:06-0500

a) The provided sample mean is "\bar{x}=38" and the sample standard deviation is "s=8."

The sample size is "n=45." The number of degrees of freedom are "df=n-1=44," and the significance level is "\alpha=0.1."

The following null and alternative hypotheses need to be tested:

"H_0: \mu=40"

"H_1: \mu\not=40"

This corresponds to two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the critical value for a two-tailed test is "t_c=1.68023."

The rejection region for this two-tailed test is "R=\{t:|t|>1.68023\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{38-40}{8/\sqrt{45}}=-1.67705"

Since it is observed that "|t|=1.67705<1.68023=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 40, at the 0.1 significance level.

Using the P-value approach: The p-value "df=44, t=-1.67705," "two-tailed," is "p=0.100634," and since "p=0.100634>0.1," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is different than 40, at the 0.1 significance level.


b) The provided sample mean is "\bar{x}=38" and the sample standard deviation is "s=8."

The sample size is "n=45." The number of degrees of freedom are "df=n-1=44," and the significance level is "\alpha=0.025."

The following null and alternative hypotheses need to be tested:

"H_0: \mu\geq40"

"H_1: \mu<40"

This corresponds to left-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the critical value for a left-tailed test is "t_c=-2.015368"

The rejection region for this left-tailed test is "R=\{t:t<-2.015368\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{38-40}{8/\sqrt{45}}=-1.67705"

Since it is observed that "t=-1.67705>-2.015368=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is les than 40, at the 0.025 significance level.

Using the P-value approach: The p-value "df=44, t=-1.67705," "left-tailed," is "p=0.0503," and since "p=0.0503>0.025," it is concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu" is less than 40, at the 0.025 significance level.